OpenStudy (anonymous):

Given the equation y − 3 = one half(x + 6) in point-slope form, identify the equation of the same line in standard form.

OpenStudy (anonymous):

walk me through it plz

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

thanks but i got someone to help

OpenStudy (anonymous):

i already wrote the equation

OpenStudy (phi):

this is the same equation, but they want it in a different form I think you should start with y = mx + b form (what you got in the last question)

OpenStudy (anonymous):

so start with the answer from the last one

OpenStudy (phi):

yes we want the form Ax + B y = C you have the equation (in y = mx +b form) \[ y = \frac{1}{2} x +6 \]

OpenStudy (anonymous):

u lost me

OpenStudy (phi):

you have \[ y − 3 = \frac{1}{2}(x + 6) \] you can change that to y = mx +b form. you get \[ y = \frac{1}{2} x+ 6 \] agreed ?

OpenStudy (anonymous):

yes we did that last time

OpenStudy (phi):

standard form means you have the x and y on the same side of the equation if any number multiplies the x (we call it A so we can talk about it), it is positive the standard form is A x + B y = C we have y = 1/2 x + 6 we want the "x term" on the other side. add -1/2 x to both sides can you do that ?

OpenStudy (anonymous):

 like this